OpenStudy (anonymous):

CHECK ANSWER. WILL MEDAL : a glass jar contains 6red, 5green, 8blue, 3yellow marbles. If a single marble is chosen at random from the jar, what is the porbability of choosing a red or green marble? My answer :: 11/20 or 23%

OpenStudy (danjs):

(red + Green) / Total

OpenStudy (danjs):

THe answer is half right

OpenStudy (anonymous):

11/20 SO WHAT IS THAT AS A PERCENT?

OpenStudy (danjs):

The total = 6 + 5 + 8 + 3 = 22

OpenStudy (anonymous):

\[ \frac {11}{20}\times \frac 55 = \frac{55}{100} = 55\% \]

OpenStudy (danjs):

11/22 *100 = your percent

OpenStudy (anonymous):

wow oops. haha

OpenStudy (anonymous):

How did you get 11/20?

OpenStudy (anonymous):

i added wrong :')

OpenStudy (danjs):

11/22 = 1/2 = a half is half of a hundred percent

OpenStudy (anonymous):

so the answer is 50%

OpenStudy (danjs):

yes